OpenStudy (anonymous):

the perimeter of the square is 208 in. Find the value of x

OpenStudy (anonymous):



OpenStudy (anonymous):

all 4 sides @Vocaloid and its 7x+3

Vocaloid (vocaloid):

hint: 4(side) = perimeter

OpenStudy (anonymous):

4+3=7?

Vocaloid (vocaloid):

side = ?

OpenStudy (anonymous):

not getting you here?

Vocaloid (vocaloid):

what is the length of one side? look at the diagram...

OpenStudy (anonymous):

(7x+3)in

Vocaloid (vocaloid):

right, and perimeter = ?

OpenStudy (anonymous):

208in

Vocaloid (vocaloid):

good, so plug side and perimeter into the equation 4*(side) = perimeter

OpenStudy (anonymous):

28+12 then =

OpenStudy (anonymous):

40

Vocaloid (vocaloid):

not quite... you're forgetting the x

Vocaloid (vocaloid):

4(7x+3) = ?

OpenStudy (anonymous):

40X

Vocaloid (vocaloid):

nope, still not quite right...

Vocaloid (vocaloid):

4(7x+3) = 4*(7x) + 4*(3) = ?

OpenStudy (anonymous):

12

Vocaloid (vocaloid):

again, you're forgetting the x... 4*(7x) = __?

OpenStudy (anonymous):

18x

OpenStudy (anonymous):

28

OpenStudy (anonymous):

x

sammixboo (sammixboo):

Well, since we know the perimeter is 208 inches, and one side is (7x +3) we can make an equation 208 = 4(7x+3) 208 is perimeter 4 is for all 4 sides (7x+3) is the length of one side. We gotta solve for x using this equation 208 = 4(7x+3), do you know how to do this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

4*7=28 4*3=12 28+12=40 but that isnt it what am i doing wrong

sammixboo (sammixboo):

Great, so our first step would be is to simplify 4(7x + 3), which is what @Vocaloid was trying to explain. Do you know what distributing is?

OpenStudy (anonymous):

yes

sammixboo (sammixboo):

You are close, just forgetting the x. Distributing is when you take the number on the outside of the parenthesis (in this case our number is 4) and multiply it by all the numbers INSIDE the parenthesis (which is 7x and 3) So we have to multiply: 4 and 7x 4 and 3 We know that 4*3 is equal to 12, so we got that down. Now for 4*7x. We have to multiply the 4 and 7, but let's not forget about the x. After you multiplied 4 and 7, throw on the x on your answer. What is 4*7?

OpenStudy (anonymous):

28